OpenStudy (anonymous):

If a=12 b=30 and c=22, find the area of triangle ABC to the nearest tenth.

OpenStudy (anonymous):



OpenStudy (anonymous):

the area is (b*h)/2

OpenStudy (anonymous):

but 12*30=360/2=180. 12*22=264/2=132. 30*22=660/2=330. But the answer is 113.1

OpenStudy (anonymous):

There are two equations that can help you tackle this question. From your other questions it looks like your working with Law of Cosines: \[\large a^2 + b^2 - 2ab *\cos(C) = c^2 \] from this formula you can find one angle given the three sides. After that you can use the sine Law to solve the remaining angles. that should break it into a reasonable question. The other method is the straight the point 'Heron's Formula' which is a means of plugging the sides directly into a formula and getting the area...tough to memorize though http://www.mathopenref.com/heronsformula.html

OpenStudy (phi):

The convention is use lower case letters to denote the lengths of sides of a triangle, so If a=12 b=30 and c=22, find the area of triangle ABC to the nearest tenth. sounds like a triangle with 3 sides of length 12, 22, and 30 Use Heron's formula to find the area of a triangle when you know the length of the 3 sides http://www.mathsisfun.com/geometry/herons-formula.html

OpenStudy (anonymous):

@phi echo echo echo